tails M y y? This question is worth 6 points. TL 115) 50 times, then 1 are mes, conducted. The ratio will be very close to

stumped on #5
Answer #1
here this is binomial with parameter n=7 and p=0.4

P(less than 60% use a gym as a routine basis )=P(less than 7*0.6 =4.2 use) =P(X<=4)

=P(X=0)+P(X=1)+P(X=2)+P(X=3)+P(X=4)

=∑x=04 (7Cx)0.4x(1−0.4)(n-x)    = 0.0280+0.1306+0.2613+0.2903+0.1935 = 0.9037